Every minute, 120 C of charge flow through this cr

Every minute, 120 C of charge flow through this cross section of the wire.The wire's current is:

A)120A

B)240A

C)2A

D)Some Other value

E)60A
Answer #1
Concepts and reason

The concepts required to solve the question is of current and charge.

Initially convert the time given in minutes into seconds form. Then use the formula of current which is charge per unit time to find the wire’s current.

Fundamentals

The charge flowing per unit time is called as current. The mathematical expression for current is,

i=dqdti = \frac{{dq}}{{dt}}

Here, i is current, q is charge, and t is time during which the flow of charge is measured.

Every minute, 120 C of charge is flowing through the cross section of the wire.

But,

1minute=(60seconds1minute)1minute1min=60sec\begin{array}{c}\\1\;{\rm{minute}} = \left( {\frac{{60\;{\rm{seconds}}}}{{1\;{\rm{minute}}}}} \right)1\;{\rm{minute}}\\\\1\;\min = 60\;\sec \\\end{array}

The mathematical expression of current is,

i=dqdti = \frac{{dq}}{{dt}}

Here, the right-hand side of the above equation states that change in charge with respect to the time gives current.

Substitute 120 C for dq and 60 s for dt.

i=120C60s=2A\begin{array}{c}\\i = \frac{{120\;{\rm{C}}}}{{60\;{\rm{s}}}}\\\\ = 2\;{\rm{A}}\\\end{array}

The current flowing through the wire is of 2 A.

Ans:

The current flowing through the wire is of 2 A.